If the complement of an angle is 22° less than one-third of its supplement, find the measure of the angle
Answer by Gogonati(855)

Let = x degree the unknown angle. Its complement ix 90-x degree and its suplement is 180-x degree.Now we write the equation as we read.
90-x=(1/3)(180-x), solving this linear equation we get:
270-3x=180-x => 2x=90 => x=45.
Answer:The unknown angle is 45 degree.